OpenStudy (anonymous):

write an equation in standard form for a line with a slope of 2 and a y intercept of 3

OpenStudy (idealist10):

y=mx+b where m=slope and b=y-intercept y=2x+3 To convert the above equation in standard form, y-2x=3 -2x+y=3 2x-y=-3